OpenStudy (butterflyhope):

What is -2 raised to the power of -3? I just need it explained. Thanks! :)

OpenStudy (cj49):

-2^-3 1/-2^3

OpenStudy (cj49):

x^-3=1/x^3

OpenStudy (anonymous):

Basically you are multiplying the number by itself by itself by the power. In this case '-3' So... It would basically be this: \[-2^{-3}= -1(-2\times-2\times-2)\] \[-1\times-8\] \[=8\]

jhonyy9 (jhonyy9):

@Matty13579 are you sure in this ?

jhonyy9 (jhonyy9):

so -2^(-3) = 1/(-2)^3 = 1/(-8) = -1/8 because -2^3 = -2*-2*-2 = 4*-2 = -8 so will be

OpenStudy (triciaal):

(-2)^-3 = (-2)^0-3 = (-2)^0 divided by (-2)^3 (-2)^0 = 1 and (-2)^3 = (-2)(-2)(-2) = -8 so it becomes 1/-8 = -1/8

OpenStudy (butterflyhope):

I'm so confused now lol xD

OpenStudy (triciaal):

when you divide numbers with the same base subtract the exponents

OpenStudy (anonymous):

@jhonyy9 Yeah, I'm sure...

OpenStudy (triciaal):

@Matty13579 sorry you made a mistake -2^-3 is not -(-2)^-3 it would be -(+2)^-3

OpenStudy (anonymous):

@triciaal Yeah, I see it now... Thanks.

OpenStudy (butterflyhope):